OpenStudy (saifoo.khan):

explain why it is necessary to have common denominators when adding or subtracting rational expressions.

OpenStudy (saifoo.khan):

Spammers will be reported.

OpenStudy (anonymous):

lol..another one!!!

hero (hero):

Because if you don't, you can't add or subtract.

OpenStudy (anonymous):

what does the word "denominator" mean?

OpenStudy (anonymous):

medal for M.H

OpenStudy (saifoo.khan):

The number below the line in a common fraction; a divisor.

OpenStudy (anonymous):

MEDALS FOR EVERYBODY!!!!

OpenStudy (anonymous):

the denominator of a rational is basically the part that's below the line

OpenStudy (anonymous):

it means "what you have" essentially. so if i have rials i can add them to rials, and if i have euros i can add them to euros, and and if i have dollars i can add them to dollars. but if i want to add euros to rials i have to convert one to the other, or else the addition makes no sense. like adding a cup of popcorn to a cup of mild

OpenStudy (anonymous):

*milk

OpenStudy (saifoo.khan):

o_O

OpenStudy (anonymous):

I HATE MILK!!

hero (hero):

Good Answer +1

OpenStudy (anonymous):

rational expressions are from mars; . . .

OpenStudy (anonymous):

why is badman spamming :-(

OpenStudy (saifoo.khan):

I never ever get help on this site.

hero (hero):

I think Satellite gave a good answer

OpenStudy (anonymous):

think about adding 1/3 to 1/5

OpenStudy (anonymous):

poor seafood!!

OpenStudy (anonymous):

adding 1/3 of a pie of pizza + 1/5 of a pie of pizza

OpenStudy (saifoo.khan):

i can't write that there. LOL

OpenStudy (anonymous):

its easier if you have a common unit or common division

OpenStudy (anonymous):

GOODNIGHT PEOPLE!!!

OpenStudy (saifoo.khan):

what i can think of is, if we divide something equally, then only we can add them correctly.

OpenStudy (anonymous):

i am happy to give a different answer, but i am not sure what kind of answer you are looking for you cannot add \[\frac{2}{x}+\frac{4}{x+5}\] without making the denominators the same. you cannot add unlike terms.

OpenStudy (saifoo.khan):

why i can't add them? that's the question

OpenStudy (saifoo.khan):

you you must have same denominators?

OpenStudy (anonymous):

we can prove it algebraically. given a/b + c/d = a*b^-1 + c*d^-1

OpenStudy (anonymous):

you can add them. just like i can add \[\frac{1}{2}+\frac{1}{3}\] or $20 + 50 cents

OpenStudy (anonymous):

BYEEEE

OpenStudy (anonymous):

now multiply by bd * (bd)^-1

OpenStudy (anonymous):

but i have to make them the same first, i.e. i have to turn them into common units.

OpenStudy (anonymous):

which is equal to 1

OpenStudy (saifoo.khan):

"common units" we can use that.

OpenStudy (anonymous):

suppose i want to add 3 feet to 4 meters. i can do it, but i have to convert one to the other first

OpenStudy (anonymous):

a/b + c/d = a*b^-1 + c*d^-1= [(bd)^-1*bd] ( a*b^-1 + c*d^-1)

OpenStudy (anonymous):

and you can assume WLOG that that gcd (b,d) = 1

OpenStudy (saifoo.khan):

can i can write: it is important to make the denominators same b/c we have to add them proportionally